OpenStudy (anonymous):

int. calculus \[\int\limits_{}^{}\sin ^{2}x \cos ^{2}x \tan ^{2} x dx\]

OpenStudy (anonymous):

Write them as sin squared x = (1/2)(1 - cos 2x) cos squared x = (1/2)(1 + cos 2x) maybe..that'll be easier for you to integrate.

OpenStudy (anonymous):

can i do that, even if sin or cos is not raised two?

hartnn (hartnn):

doesn't that simplify to sin^4 x when tan is written as sin/cos ?

OpenStudy (anonymous):

Write tan^2 x = sin^2 x / cos^2 x

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i did that before

OpenStudy (anonymous):

\[\large{\int { sin^4x dx}}\] Can you do this?

OpenStudy (anonymous):

then i lost any of ideas after that

OpenStudy (anonymous):

OK no problem, where are you lost?

hartnn (hartnn):

can u integrate sin^2 x ?

OpenStudy (anonymous):

after i got the \[\int\limits_{?}^{?}\sin ^{4}xdx\]

OpenStudy (anonymous):

Well if integration of sin x dx= -cosx + C then that of sin^4 x dx = ?

OpenStudy (anonymous):

-cos^4x + C ??

hartnn (hartnn):

write sin^4 x = (sin^2 x)(1-cos^2 x)

hartnn (hartnn):

no, its not that direct

OpenStudy (anonymous):

then distribute?

hartnn (hartnn):

= (sin^2 x)(1-cos^2 x) = sin^2 x - sin^2x cos^2 x yes

hartnn (hartnn):

= sin^2 x - sin^2x cos^2 x = sin^2 x - (1/4) (sin 2x)^2

OpenStudy (anonymous):

how did you get (1/4)(sin 2x)^2?

hartnn (hartnn):

sin 2x = 2 sin x cos x to get sin^2x cos^2 x , i squared the above .....

hartnn (hartnn):

got that ? ^ and can u integrate sin^2 x - (1/4) (sin 2x)^2

OpenStudy (anonymous):

can you show me the whole process of that sin 2x = 2 sin x cos x

hartnn (hartnn):

u want me to prove 2sin x cos x = sin 2x or u wanted this : sin 2x = 2 sin x cos x (sin 2x)^2 = (2 sin x cos x)^2 = 4 sin^2x cos^2 x so sin^2 x cos^2 x = (1/4) (sin 2x)^2

hartnn (hartnn):

so if u can integrate sin^2 x you can also integrate sin^2 x - (1/4) (sin 2x)^2